Find the center and radius of the circle given by the equation
x^2+y^2+4x+4y−28=0

Answer by stanbon(75887)

Find the center and radius of the circle given by the equation
x^2+y^2+4x+4y−28=0
-----
Complete the square on the x-terms and on the y-terms:
x^2 + 4x + 4 + y^2 + 4y + 4 = 28 + 4 + 4
------------
(x+2)^2 + (y+2)^2 = 34
----
center: (-2,-2)
radius: sqrt(34)
======================
Cheers,
stan H.